remi2019:

What is the value of y in the equation 2(3y + 6 + 3) = 196 − 16? 27 29 30 33

jhonyy9:

like a first step distribute parentheses

remi2019:

36?

jhonyy9:

what is this 36 ?

sxdsouls:

Step 1: Simplify both sides of the equation. 2(3y+6+3)=196−16 (2)(3y)+(2)(6)+(2)(3)=196+−16(Distribute) 6y+12+6=196+−16 (6y)+(12+6)=(196+−16)(Combine Like Terms) 6y+18=180 6y+18=180 Step 2: Subtract 18 from both sides. 6y+18−18=180−18 6y=162 Step 3: Divide both sides by 6. 6y 6 = 162 6 y=27 sorry if this is direct, idk- dlfdjskljflkdsj. or do i delete- cdlfjkd

remi2019:

thx

jhonyy9:

good job @sxdsouls

remi2019:

sxdouls that helped and i got it right

sxdsouls:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 good job @sxdsouls \(\color{#0cbb34}{\text{End of Quote}}\) Ofc, glad to help users.

sxdsouls:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @remi2019 sxdouls that helped and i got it right \(\color{#0cbb34}{\text{End of Quote}}\) glad you did <3

jhonyy9:

yw

caydenisdead:

27